Dear Maths expert,
Please I need some help with the following mini-case:
Below are the results of a survey:
Major
Male - Female
Accounting 100 (M) - 100 (F)
Management 150 (M) - 50(F)
Finance 50 (M)- 50(F)
Total 300 (M)- 200(F)
A) What is the probability of selecting a female student?
B) What is the probability of selecting a finance or accounting major?
C) What is the probability of selecting an accounting major, given that the person selected is a male?
Thanks very much in advance.
Gerald

Answer by funmath(2933)

Hi Gerald,
Way to butter us up with the whole "math expert" thing!
:
Below are the results of a survey:
:
Major
Male - Female
Accounting 100 (M) - 100 (F)
Management 150 (M) - 50(F)
Finance 50 (M)- 50(F)
Total 300 (M)- 200(F)
A) What is the probability of selecting a female student?
The total number of Females over the total numeber of students is:
P=200/(200+300)=200/500=2/5
:
B) What is the probability of selecting a finance or accounting major?
Total Number of finance students over total number of students + total number of accounting majors over the total number of students is:
P=(100+100)/500+(150+50)/500=200/500+200/500=400/500=4/5
:
C) What is the probability of selecting an accounting major, given that the person selected is a male?
The total number of accounting majors that are male over the total number of males is:
p=100/300=1/3
:
Happy Calculating, Gerald!!!